OpenStudy (anonymous):

X + Y + Z = 36 4X + 2Y + Z = 71 6X + 5Y + 4Z = 86 HELP!! System of equations.

OpenStudy (blurbendy):

have you tried the elimination method?

OpenStudy (anonymous):

I'm really honestly confused with this...

OpenStudy (blurbendy):

the elimination method?

OpenStudy (anonymous):

Yes, how do I do that?!

OpenStudy (blurbendy):

it's fun. first pick a variable you want to KILL

OpenStudy (anonymous):

I'm taking online classes but it doesn't explain how to do it...

OpenStudy (blurbendy):

i will show you. choose a variable you want to eliminate. ANY ONE.

OpenStudy (kainui):

KILL THEM ALL!

OpenStudy (anonymous):

Ahh, thank's! x?

OpenStudy (blurbendy):

DIE!

OpenStudy (anonymous):

tired matrix method?

OpenStudy (blurbendy):

Let's kill 4x, it will be easier. here's why: in order to eliminate 4x, all we have to do is multiply the top equation by -4. if we do that, we're left with -4x -4Y -4Z = -144 + 4x + 2Y +Z = 71 Now add these two equations

OpenStudy (anonymous):

Will do! So, together I got.... x-2y-3z=-73. Not sure if it's correct.

OpenStudy (anonymous):

Is that how you do it?

OpenStudy (blurbendy):

close, -4x + 4x = 0, s the x variable is gone now

OpenStudy (anonymous):

Oh, well that wasn't close enough. :(

OpenStudy (anonymous):